### v1

# thanks: Equal contributionthanks: Equal contribution

Abstract
Interpreting multi-messenger signals from neutron stars and black holes requires reliable general-relativistic magnetohydrodynamics (GRMHD) simulations across rapidly evolving high-performance-computing platforms, yet key algorithms are routinely rewritten within infrastructure-specific numerical-relativity codes, hindering verification and reuse. We present the General Relativistic Hydrodynamics Library (GRHayL), a modular, infrastructure-agnostic GR(M)HD library providing conservative-to-primitive recovery, reconstruction, flux/source and induction operators, equations of state, and neutrino leakage through an intuitive interface. GRHayL refactors and extends the mature IllinoisGRMHD code into reusable pointwise and stencil-wise kernels, enabling rapid development and cross-code validation in diverse frameworks, while easing adoption of new microphysics and future accelerators. We implement the same kernels in the Einstein Toolkit (Carpet and CarpetX) and BlackHoles@Home, demonstrating portability with minimal duplication. Validation combines continuous-integration unit tests with cross-infrastructure comparisons of analytic GRMHD Riemann problems, dynamical Tolman–Oppenheimer–Volkoff evolutions, and binary neutron-star mergers, showing comparable or improved behavior over legacy IllinoisGRMHD and established Einstein Toolkit codes.
I Introduction
Over the past decade, gravitational wave (GW) astronomy has experienced rapid growth in scientific output, starting with the first direct detection of GWs [GW150914] and now encompassing over 200 compact binary coalescence detections [GW151226, GW170104, GW170608, GW170814, GW170817, GW190412, GW190521, GW190521_analysis, GW190814, GW200105, LIGO_catalog1, LIGO_catalog2, LIGO_catalog3, LIGO_catalog4, LIGO_catalog5]. Among these first observations, perhaps the most impactful was GW170817/GRB170817A/AT2017gfo [GW170817], the only multi-messenger binary neutron star (BNS) merger detection to date. The GW signal was accompanied by observations of gamma rays and months-long emission throughout much of the rest of the electromagnetic (EM) spectrum [GW170817_GRB, Goldstein_2017, Tanvir_2017]. This event is expected to be just the first in GW-anchored multi-messenger astronomy, which combines data from GW signals with data from coincident EM and, for sufficiently nearby sources, possibly neutrino detections to investigate some of the most extreme events in the cosmos. Looking forward, iterative improvements to the Advanced LIGO detectors, as well as future ground-based [CosmicExplorer1, CosmicExplorer2, punturo2010einstein] and space-based [LISA_2023, Luo_2016] GW detectors, will collectively provide remarkably improved detection capabilities, triggering neutrino and electromagnetic follow-ups across the spectrum.
Beyond BNS mergers, multi-messenger observations of black hole-neutron star (BHNS) systems, circumbinary disks, core-collapse supernovae, and other compact-object environments are poised to greatly expand our understanding of high-energy astrophysical systems. However, these insights are gained by comparing observations with theoretical models, which in turn will be limited by how faithfully they track the underlying physics.
Generating self-consistent models of important multi-messenger sources is exceedingly challenging, as it in principle requires fully solving the equations of general relativity (GR) for the spacetime, combined with a magnetohydrodynamic (MHD) treatment of turbulent magnetized plasmas, nuclear equations of state (EoSs), neutrino and EM radiation transport, r-process nucleosynthesis via nuclear reaction networks, and detailed EM ray tracing. Fully self-consistent modeling of all of these processes in a single three-dimensional simulation exceeds the computational resources currently available to the community, not to mention that the relevant length and time scales span several orders of magnitude, making even one such calculation prohibitively expensive. As NR codes are extended to incorporate additional microphysics and radiation effects, the underlying frameworks have historically counterbalanced the growing computational cost by taking advantage of advances in central processing unit (CPU) architectures. However, just as the revolution in multi-messenger astronomy has greatly increased the importance of improving and extending the physical realism of NR models, the rapid shift towards heterogeneous high-performance computing (HPC) systems and reliance on accelerators like graphics processing units (GPUs)—heavily driven by the advent of artificial intelligence and machine learning—now require that code frameworks be upgraded to take full advantage of state-of-the-art HPC resources.
Transforming existing CPU-tuned NR frameworks to be performant and scalable on heterogeneous CPU/GPU HPC clusters typically requires a substantial redesign of the underlying software infrastructure. As a result, significant effort has been expended in recent years to migrate existing simulation software to computational frameworks that explicitly support heterogeneous architectures and offer performance portability across devices. Many groups are working towards extending their in-house infrastructures [Ishi_2019] and/or developing entirely new ones [ET_2023_11] that build on performance-portable programming models or modern AMR infrastructures in order to fully exploit current and future HPC resources.
Two notable examples of technologies enabling this transition are Kokkos [kokkos] and AMReX [amrex], which play complementary roles. Kokkos is a general-purpose performance-portability library that provides an abstraction layer for expressing parallelism and memory access patterns in a way that can be mapped efficiently onto diverse heterogeneous HPC architectures. In the NR community, Kokkos has been used to extend Athena++ [Stone_2020, GR_Athena] to be GPU-capable. The Parthenon framework [Grete:2023] further develops the original Athena++ code into a more general framework for multiphysics code development on top of Kokkos. Several other projects connected to the Los Alamos National Laboratory also provide open-source components built on Kokkos that are directly relevant to NR, such as the portability abstraction library Ports-of-Call [PortsOfCall], the tabulated data interpolation library Spiner [miller2022notquite, Miller2022], and the EoS library Singularity EOS [miller2024singularity].
In parallel, AMReX provides a general framework for implementing block-structured adaptive mesh refinement (AMR) codes on heterogeneous HPC systems. It supplies the mesh-management, load-balancing, and parallelization infrastructure needed to build scalable AMR applications on modern CPU/GPU clusters, and has been used by several groups to develop NR and NR-adjacent codes for these systems. The Einstein Toolkit [ET_2023_11], primarily built on the Cactus Framework [Cactus] and the Carpet AMR driver, is one example of an existing code base undergoing such a modernization. Recently, the new AMReX-based CarpetX code was developed and released in the Einstein Toolkit as a GPU-capable replacement for Carpet. Another notable example is GRChombo [Clough_2015, Andrade2021], whose developers are in the process of creating GRTeclyn and porting the science code to the new infrastructure.
As all these efforts show, new infrastructures are required to fully leverage current and future developments in HPC and numerical techniques. Thus, in order to keep pace with these changes, there is a clear and urgent need to minimize the time and effort required to migrate NR codes between infrastructures. The current paradigm often requires significant code rewrites or refactoring, which is undesirable, as most GRMHD algorithms are independent of the infrastructure.
Several groups have released code packages that partially address this issue. For instance, several open-source codes provide functions for conservative-to-primitive recovery schemes [Noble:2005gf, Siegel:2017jug] and interpolators for finite-temperature EoS tables [OConnor:2009iuz, Schneider:2017tfi]. While these codes require some work to incorporate into a new code, they do provide core algorithms in a reasonably self-contained form. Additionally, several libraries have been developed to provide implementations of EoS [miller2022notquite, Miller2022, miller2024singularity], conservative-to-primitive recovery schemes [Kastaun_2021, Roland_2024], and neutrino physics [OConnor:2014sgn, Cheong:2023fgh].
These groups have demonstrated both the success of and demand for infrastructure- and hardware-agnostic implementations of the core GRMHD algorithms needed for NR simulations. To more comprehensively address the needs of GRMHD simulations, we introduce the General Relativistic Hydrodynamics Library (GRHayL). This library contains all the core elements required for a GRMHD simulation and represents a refactoring and extension of the IllinoisGRMHD code [Duez_IGM, Etienne_IGM]—including the addition of tabulated EoS and neutrino leakage support in the forked code by Werneck et al. [Werneck_IGM]—into a modular, infrastructure-agnostic library. GRHayL provides nearly all of the code required to develop an IllinoisGRMHD-like code via pointwise or stencil-wise library functions, allowing for quicker code development, easier cross-code and cross-infrastructure collaboration, and cleaner implementations of GRMHD codes. There is also ongoing work to add full GPU support.
This paper is organized as follows. Section II discusses the individual code modules within the library and provides an overview of the available functionality. Section III lists all the functional changes and improvements we have made to the IllinoisGRMHD code while developing the library. Section IV describes the codes we have developed using GRHayL as well as the infrastructures themselves, in preparation for Section V, which compares data from the original IllinoisGRMHD code and various new implementations using GRHayL. Finally, in Section VI we summarize the library features and discuss planned extensions of GRHayL’s capabilities and implementations in additional infrastructures. In addition, we include two appendices. Appendix A briefly describes the GRMHD evolution equations and our choice of variables used for evolving them. Appendix B details the automated, function-level continuous integration testing that we perform to validate the library for each change made to the code repository.
II Library Features
The main purpose of GRHayL is to provide an infrastructure-agnostic, modular, and easily extensible GRMHD library. A key step in achieving modularity is the identification of the code’s core algorithms, which in the context of GRMHD include primitive recovery routines, primitive variable reconstruction, and computation of the flux and source terms of the GRMHD equations. It is important to note that the EoS is unique in that many parts of the code depend on it directly, so special care is required to expose it through a clean, modular interface.
As part of developing GRHayL, several improvements have been made to the original IllinoisGRMHD code. To clarify which version we are referencing, we introduce the following terminology. The original IllinoisGRMHD and the Werneck et al. fork are collectively referred to as origIGM, and these are the versions of IllinoisGRMHD that have been incorporated into GRHayL. We refer to the new GRHayL-based code as IllinoisGRMHD.
Regarding extensibility, GRHayL was designed to facilitate the seamless integration of new modules (or “gems”), such as neutrino physics, without disturbing the existing framework. Notably, the separation is such that if one were to replace GRHayL’s conservative-to-primitive routines with an alternative, the rest of the code would remain unaffected. This is a fundamental aspect of GRHayL’s design, offering the flexibility for users to either extend GRHayL with their own custom routines or directly incorporate new routines into their code while leveraging the library’s full capabilities.
The core part of the library—the “chalice” of GRHayL—is the glue that ties all the separate modules together. This must be as minimal as possible to meet the aforementioned requirements of the library. GRHayL’s core contains definitions of C structs and simple helper functions for packing and unpacking data to and from these structs, as well as performing basic operations dependent purely on struct quantities (e.g., computing $g_{\mu\nu}$ and $g^{\mu\nu}$ from the ADM variables).
Affixed to this chalice are the various gems that implement specific features. The currently provided gems are Atmosphere, Con2Prim, EOS, Flux_Source, Induction, Neutrinos, and Reconstruction. The Atmosphere gem is quite simple, as it currently only contains the constant-density atmosphere prescription used by origIGM. Other prescriptions, such as an atmosphere where hydrodynamic quantities follow a radial falloff, will be added in the future. We discuss the other gems in the rest of this section.
II.1 EOS Gem
The EOS gem provides EoS functionality within GRHayL. It currently supports hybrid piecewise polytropic EoSs and includes infrastructure for fully tabulated EoSs. Tabulated EoS support is presently under development and is not yet intended for production simulations. Many other gems depend on functions provided by the EOS gem, but this dependency has been abstracted to allow users to implement their own EoS driver without changing the other gems. The EoS information is conveyed primarily via an EOS struct, in addition to several function pointers that are set during initialization. Users are free to swap the EOS functions with their own by setting these function pointers. Further, other EoS types or implementations can be easily added to the gem. For example, providing support for the Singularity EOS library would simply require defining the proper wrapper functions that convert the expected arguments for the GRHayL functions and making the appropriate call to Singularity EOS.
A hybrid EoS [1993A&A...268..360J] assumes that the pressure and specific internal energy can be split into cold and thermal components, i.e.,
$P(\rho)$ $=P_{\rm cold}(\rho)+P_{\mathrm{th}}(\rho)\;,$
$\epsilon(\rho)$ $=\epsilon_{\mathrm{cold}}(\rho)+\epsilon_{\mathrm{th}}(\rho)\;.$
In GRHayL, the cold pressure is represented by a piecewise polytrope (see, e.g., [Read:2008iy]),
$P_{\rm cold}(\rho)=\left\{\begin{matrix}[l]K_{0}\rho^{\Gamma_{0}}&,&\text{if }\rho\leq\rho_{0}\;,\\ K_{1}\rho^{\Gamma_{1}}&,&\text{if }\rho_{0}<\rho\leq\rho_{1}\;,\\ \hskip 11.38092pt\vdots&&\hskip 28.45274pt\vdots\\ K_{n}\rho^{\Gamma_{n}}&,&\text{if }\rho>\rho_{n-1}\;,\end{matrix}\right$.
where $\bigl(K_{0},\{\Gamma_{i}\},\{\rho_{i}\}\bigr)$ are user-provided parameters. The cold specific internal energy is determined via the integral
$\epsilon_{\mathrm{cold}}(\rho)=\int d\rho^{\prime}\,\frac{P_{\rm cold}(\rho^{\prime})}{\rho^{\prime 2}}\;,$
assuming that $\lim\limits_{\rho\to 0}\epsilon_{\mathrm{cold}}=0$. The thermal component of the EoS is determined using
$P_{\mathrm{th}}=\bigl(\Gamma_{\mathrm{th}}-1\bigr)\rho\epsilon_{\mathrm{th}}=\bigl(\Gamma_{\mathrm{th}}-1\bigr)\rho\bigl(\epsilon-\epsilon_{\mathrm{cold}}\bigr)\;,$
where $\Gamma_{\mathrm{th}}$ is a constant parameter that determines the conversion efficiency of kinetic to thermal energy at shocks. The EOS gem additionally supports a special case of the hybrid EoS called the simple EoS. [^2] This is often referred to as the “ideal fluid”, “ideal gas”, or “Gamma-law” EoS. In this case, the EoS contains a single piece with $\Gamma_{0}=\Gamma_{\mathrm{th}}$, resulting in $P=(\Gamma_{0}-1)\rho\epsilon$.
II.2 Con2Prim Gem
The Con2Prim gem provides several methods for recovering the primitive variables $\bm{P}$ from the conservative variables $\bm{C}$ (defined in Eqs. 12 and 16, respectively). Because of the nonlinearity of the equations, a root-finding method such as Newton–Raphson is required. In addition, this gem provides several helper functions to compute $\bm{C}(\bm{P})$, to produce initial guesses for the primitives, and to enforce bounds on $\bm{C}$ [Faber1, Faber2] and $\bm{P}$.
The gem includes several primitive recovery routines, including some by Noble et al. [Noble:2005gf, Noble:2008tm], Font et al. [Font_2000, Etienne_2012_3], Palenzuela et al. [Palenzuela_2015], and ongoing efforts to support Newman & Hamlin [Newman_2014] (see also [Werneck_IGM] for an entropy variation of the last two). Routines are named based on the first author of their originating paper and the dimensionality of the method (e.g., Noble2D, Palenzuela1D, etc.). The gem also provides several useful diagnostics, including the number of iterations a routine took to converge to the solution and whether or not backup routines were used.
The routines are adapted from origIGM and from [Siegel:2017jug], noting that origIGM originally adapted the Noble et al. routines from HARM3D [Noble:2005gf, Noble:2008tm] and HARM3D+NUC [Murguia-Berthier:2021tnt]. We summarize all available primitive recovery routines currently implemented in GRHayL in Table 1.
Table 1: List of provided conservative-to-primitive routines and their compatibility with the different EoS options.
Method Equation of State
Simple Hybrid
Noble2D [Noble:2005gf]
Noble1D [Noble:2005gf]
Noble1D_entropy [Noble:2008tm]
Font1D [Font_2000, Etienne_2012_3]
Palenzuela1D [Palenzuela_2015]
Palenzuela1D_entropy [Werneck_IGM]
II.3 Flux_Source Gem
The Flux_Source gem provides functions to compute the flux and source terms for the right-hand sides (RHSs) of the evolved GRMHD variables as described in Appendix A. These functions calculate the flux contributions to the RHSs using a finite-volume method, and the needed reconstructions can be performed using the Reconstruction gem. Flux_Source uses NRPy [Ruchlin:2018] to automatically generate optimized C code from the basic flux and source equations. The Flux_Source gem currently only provides the Harten, Lax, & van Leer (HLL) [Harten_1983] approximate Riemann solver, which has the form
$F^{\mathrm{HLL}}=\frac{c^{-}F_{r}+c^{+}F_{l}-c^{+}c^{-}\left(U_{r}-U_{l}\right)}{c^{+}+c^{-}}\;.$ (1)
Here $c^{-}$ is the minimum characteristic speed, $c^{+}$ is the maximum characteristic speed, $F_{r,l}$ are the hydrodynamic fluxes defined in Eq. (20), and $U_{r,l}$ are the reconstructed values for the conserved variable $U$. The $r,l$ subscripts indicate fluid and spacetime quantities reconstructed at the right and left faces along some $i$-th flux direction. To compute the characteristic speeds, we follow the prescription given by [HARM], which approximates the general GRMHD dispersion relation as a quadratic in the wave speed. The wave speeds are then given by evaluating the quadratic equations
$c^{+}$ $=\max\left(\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}\right)\;,$
$c^{-}$ $=\min\left(\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}\right)\;,$
with coefficients given by
$a$ $=\left(1-v_{0}^{2}\right)\left(u^{0}\right)^{2}-v_{0}^{2}g^{00}\;,$
$b$ $=2v_{0}^{2}g^{i0}-2u^{i}u^{0}\left(1-v^{2}_{0}\right)\;,$
$c$ $=\left(1-v_{0}^{2}\right)\left(u^{i}\right)^{2}-v_{0}^{2}g^{ii}\;,$
where $v_{\mathrm{A}}$ is the Alfvén speed and $c_{\mathrm{s}}$ the sound speed. The variable $v_{0}$ is defined as
$v_{0}^{2}\equiv\frac{\omega^{2}}{k^{2}}=v_{\mathrm{A}}^{2}+c_{\mathrm{s}}^{2}\left(1-v_{\mathrm{A}}^{2}\right)\;,$
where $\omega$ is the wave angular frequency and $k$ is the wave number. Since the characteristic speeds are computed at each cell interface from reconstructed variables, GRHayL returns final characteristic speeds
$c^{+}=\max\left(0,\ c^{+}_{l},\ c^{+}_{r}\right)\text{ and }c^{-}=-\min\left(0,\ c^{-}_{l},\ c^{-}_{r}\right)\;.$
We note that the characteristic speeds obtained from these equations are slightly overestimated, resulting in a modest increase in numerical diffusion and consequently in robustness.
Conforming to GRHayL’s design principles, this gem computes the fluxes and source terms in a pointwise fashion. Reconstructed primitives, face-interpolated values of metric quantities, and spatial derivatives of metric quantities are taken as inputs, decoupling the gem from the method used to compute them. The functions within the gem are naturally divided into three subcategories: characteristic speeds, fluxes, and source terms. All of these use EoS functions to compute the specific enthalpy, with the characteristic speed functions also computing the sound speed. Further, the characteristic speeds and fluxes are each split into direction-dependent versions. Finally, we provide variations of the flux functions to handle the following four possible evolution options: hybrid EoS, hybrid EoS with entropy evolution, tabulated EoS, and tabulated EoS with entropy evolution.
II.4 Induction Gem
To preserve the divergence-free nature of the magnetic field, instead of directly evolving $\tilde{B}^{i}$ using the flux terms above for the induction equation, we evolve the vector potential $A_{k}$ using a staggered constrained-transport scheme [Balsara:1999, Etienne:2012, DelZanna_2003]. This guarantees that $\tilde{B}^{i}=\sqrt{\gamma}B^{i}$ remains divergence-free because, before each calculation of the evolution equations we compute
$\tilde{B}^{i}=\epsilon^{ijk}\partial_{j}A_{k}\;,$ (2)
where $\epsilon^{ijk}=\epsilon_{ijk}$ is the totally antisymmetric Levi-Civita symbol with $\epsilon^{xyz}=1=\epsilon_{xyz}$.
The Induction gem provides functions to facilitate the computation of the RHSs of the induction evolution equations in the Lorenz gauge. These are given by
$\partial_{t}A_{i}$ $=\epsilon_{ijk}v^{j}\tilde{B}^{k}-\partial_{i}(\alpha\Phi-\beta^{j}A_{j})\;,$ (3)
$\partial_{t}\tilde{\Phi}$ $=-\partial_{i}(\alpha\sqrt{\gamma}A^{i}-\beta^{i}\tilde{\Phi})-\lambda\alpha\tilde{\Phi}\;$, (4)
where the second terms in both equations are gauge terms. Here, $\Phi$ is the scalar potential, $\tilde{\Phi}=\sqrt{\gamma}\Phi$, and $\lambda$ is the damping factor of the ‘generalized Lorenz gauge condition.’
To illustrate how the non-gauge terms in Eq. (3) are computed, let $\mathcal{E}_{i}\equiv-\epsilon_{ijk}v^{j}\tilde{B}^{k}$. Focusing on, e.g., the $z$-component and dropping the gauge terms, we find
$\partial_{t}A_{z}^{\left(i+\frac{1}{2},j+\frac{1}{2},k\right)}=-\mathcal{E}_{z}^{\left(i+\frac{1}{2},j+\frac{1}{2},k\right)}.$
We then use a two-dimensional HLL approximate Riemann solver to determine $\mathcal{E}_{i}$ so that it is consistent with the staggering of the vector potential. As Eq. (3) suggests, this is most naturally done using the densitized magnetic field $\tilde{B}^{i}=\sqrt{\gamma}B^{i}$. Continuing to focus on the $z$-component, we have
$\left(\mathcal{E}_{z}\right)^{\rm HLL}=\frac{c_{xy}^{++}\mathcal{E}_{z}^{\rm LL}+c_{xy}^{+-}\mathcal{E}_{z}^{\rm LR}+c_{xy}^{-+}\mathcal{E}_{z}^{\rm RL}+c_{xy}^{--}\mathcal{E}_{z}^{\rm RR}}{(c_{x}^{+}+c_{x}^{-})(c_{y}^{+}+c_{y}^{-})}$
$\quad+\frac{c_{xx}^{+-}}{c_{x}^{+}+c_{x}^{-}}\bigl(\tilde{B}^{y}_{\rm R}-\tilde{B}^{y}_{\rm L}\bigr)-\frac{c_{yy}^{+-}}{c_{y}^{+}+c_{y}^{-}}\bigl(\tilde{B}^{x}_{\rm R}-\tilde{B}^{x}_{\rm L}\bigr),$ (5)
where $c_{ab}^{mn}\equiv c_{a}^{m}c_{b}^{n}$, $\mathcal{E}^{\rm LR}_{z}$ indicates a quadrant reconstruction that is left in $x$ and right in $y$; similarly $\rm LL,RL,RR$} label the other $xy$-quadrants. The subscripts $\rm L/R$ on $\tilde{B}^{y}$ denote left/right with respect to the $x$-direction, while $\rm L/R$ on $\tilde{B}^{x}$ denote left/right with respect to the $y$-direction. The formula for $\left(\mathcal{E}_{x}\right)^{\rm HLL}$ can be obtained from Eq. (5) via cyclic permutation of the indices, $(z,x,y)\to(x,y,z)$, and for $\left(\mathcal{E}_{y}\right)^{\rm HLL}$ via double cyclic permutation, $(z,x,y)\to(y,z,x)$.
Given the cyclic nature of the expression, the directional nature of the curl in our implementation is handled by simply providing the correct components in the input data. The required inputs are the velocities and magnetic fields reconstructed to the location of $A_{i}$, as well as the characteristic speeds $c^{\pm}$. Functions to compute the characteristic speeds are provided in the Flux_Source gem, since the hydrodynamic fluxes also require $c^{\pm}$.
Having specified how the non-gauge flux term is computed, we now turn to the interpolation needed for the gauge-related terms and the scalar potential. As with the hydrodynamic RHS, $\tilde{\Phi}^{\mathrm{RHS}}$ and the gauge contribution to $A_{i}^{\mathrm{RHS}}$ require staggered variable quantities. However, we can use interpolation instead of reconstruction methods because gauge variables are not as sensitive to shocks. The Induction gem provides interpolator functions to facilitate the calculation of these terms.
The primary difficulty with these quantities is that they are all sampled at different cell locations, which complicates the interpolation scheme. The hydrodynamic quantities are cell-centered, and we denote these gridpoints as $(i,j,k)$, $(i+1,j,k)$, etc. The vector potential components are edge-centered—e.g., $A_{x}$ is evaluated at the point (i,j+1/2,k+1/2)(i,j+1/2,k+1/2)—and the scalar potential is vertex-centered. This staggering is summarized in Table 2.
Table 2: Staggering of vector and scalar potentials relative to the cell-centered hydrodynamic quantities at $(i,j,k)$.
Variable $\bm{x}$ $\bm{y}$ $\bm{z}$
$A_{x}$ $i$ $j+\frac{1}{2}$ $k+\frac{1}{2}$
$A_{y}$ $i+\frac{1}{2}$ $j$ $k+\frac{1}{2}$
$A_{z}$ $i+\frac{1}{2}$ $j+\frac{1}{2}$ $k$
$\tilde{\Phi}$ $i+\frac{1}{2}$ $j+\frac{1}{2}$ $k+\frac{1}{2}$
For $\tilde{\Phi}^{\mathrm{RHS}}$ and the remaining term in $A_{i}^{\mathrm{RHS}}$, we need to interpolate the vector potential to all the various staggered positions. We also need to interpolate the lapse and the shift. Finally, we need the metric to raise the index of $A_{i}$.
This dependency on the metric leads to several choices in how to compute the RHS. The original origIGM code uses the BSSN metric, but it can be more convenient in some codes to use the ADM metric. Additionally, different infrastructures assume the spacetime has different centerings. For example, the Carpet driver in the Einstein Toolkit assumes all variables have the same centering, so the spacetime is implicitly cell-centered (i.e., it is on the same grid as the hydrodynamic variables), while CarpetX supports different grid types and has vertex-centered spacetime variables but cell-centered hydrodynamic variables. Thus, we currently provide interpolators using the ADM metric with cell- and vertex-centered variables, and the BSSN metric with cell-centered variables.
These choices only affect spacetime variable inputs, but the outputs are the same, returning all the interpolated quantities necessary to compute the remaining RHS terms. For the gauge terms in Eq. (3), the interpolated quantity is all that is needed, and the user can then compute the spatial derivative to the desired order using a standard finite-difference algorithm. We also provide all the interpolated quantities to compute the RHS in Eq. (4). In addition, we provide a function for computing the RHS of Eq. (4) that automatically implements upwinding in the spatial derivative of $\beta^{i}\tilde{\Phi}$, consistent with the usual BSSN shift upwinding.
II.5 Reconstruction Gem
To use the HLL [Harten_1983] flux solver described in Sec. II.3, the primitive variables must be interpolated to cell faces without introducing oscillatory behavior. The Reconstruction gem provides several methods to perform this interpolation. The library offers minmod, monotonized-central, and superbee reconstruction, which apply total variation diminishing (TVD) linear interpolations [TVD] to the primitives. We refer to these methods collectively as the piecewise linear methods (PLM), noting that they are at best second-order accurate for smooth underlying solutions and revert to first-order at shocks.
GRHayL also provides the piecewise parabolic method (PPM) [PPM1, PPM2], which is third-order accurate in smooth regions and first-order accurate at extrema. We note that, other than PPM, all the reconstruction functions have the same argument lists, making switching between methods very simple. The PPM scheme requires additional inputs for its special handling of shocks and thus does not use the same interface. GRHayL initializes all PPM parameters to standard values by default, but the user can tune these parameters to achieve different behavior.
II.6 Neutrinos Gem
GRHayL adopts what is perhaps the most popular approach for modeling neutrino physics in GRMHD simulations: a leakage scheme [vanRiper:1981mko, Ruffert:1995fs, Rosswog:2002rt, Rosswog:2003rv, Sekiguchi:2010ep, Sekiguchi:2011zd]. In this type of scheme, experimental data are used to parametrize analytic formulas for the neutrino emission and cooling rates in terms of the optical depths and opacities, resulting in a computationally inexpensive algorithm [OConnor:2009iuz, Ott:2012kr, Neilsen:2014hha, Radice:2016dwd, Siegel:2017jug, Endrizzi:2019trv, Murguia-Berthier:2021tnt, Werneck_IGM]. However, this scheme neglects the absorption of neutrinos and fails to account for heating and lepton number changes within hot ejecta from compact object mergers [Just:2021cls]. To account for these effects, more sophisticated techniques, like radiation transport with an M1 closure [1981MNRAS.194..439T, Shibata:2011kx] (see also [Richers:2020ntq, Radice:2021jtw]) or Monte Carlo methods [Miller:2019gig, Miller:2019dpt, Foucart:2020qjb, Foucart:2021mcb, Foucart:2021ikp], are necessary.
The implementation of the leakage scheme in GRHayL is provided by NRPyLeakage [Werneck_IGM]. [^3] We note that the neutrino source term for $\tau$ shown in Eq. (21) of [Werneck_IGM] is incorrect, as it is missing a factor of the lapse function. Our earlier implementation used the form $\sqrt{\gamma}\alpha u^{0}\mathcal{Q}=\sqrt{\gamma}W\mathcal{Q}$, but the correct contribution is $\alpha\sqrt{\gamma}W\mathcal{Q}$. This typo was also present in our implementation of the source term, but has since been fixed. Following [Ruffert:1995fs, Galeazzi:2013mia, Siegel:2017jug, Murguia-Berthier:2021tnt, Burrows:2004vq], we consider neutrino production from $\beta$-processes (electron/positron absorption by protons/neutrons), electron-positron pair annihilation, transverse plasmon decay, and nucleon-nucleon bremsstrahlung. The inverse reactions contribute to the total neutrino transport opacities, with inverse $\beta$-processes providing the dominant contribution to the optical depths of electron neutrinos and neutral-current scattering off neutrons the dominant contribution to the optical depths of heavy-lepton neutrinos.
For computing the optical depths, we adopt the general-purpose method of Nielsen et al. [Neilsen:2014hha] (see also [Siegel:2017jug, Murguia-Berthier:2021tnt, Werneck_IGM]), in which neutrinos leave the system by following the path of least resistance. Refer to Werneck et al. [Werneck_IGM] for more details on NRPyLeakage.
Given that the EOS gem’s tabulated EoS support is currently experimental (and not covered in this paper), it is not recommended as the basis for production simulations. Nevertheless, the Neutrinos gem remains suitable for production use when interfaced with an external EoS implementation or infrastructure that provides the needed thermodynamic quantities.
III Improvements over origIGM
Thus far, we have primarily discussed the conversion of the core origIGM routines into an infrastructure-agnostic library. In addition to simplifying and improving some of its core algorithms, we also found many ways to improve the code that ties these algorithms together when implementing IllinoisGRMHD. As a reminder, we use origIGM to refer collectively to the original IllinoisGRMHD thorn and the enhanced fork with neutrino leakage and experimental tabulated EoS support, while IllinoisGRMHD refers to the new version of the thorn that uses GRHayL. In addition, we introduce the new GRHayLHD thorn, which provides a purely hydrodynamic evolution code using GRHayL.
III.1 Code Simplification
Over time, repeated inheritance of code has led to the build-up of technical debt within origIGM. While refactoring the code to write GRHayL, we found several places where the code could be simplified or condensed, with most of these improvements affecting the conservative-to-primitive routine and its surrounding functions.
One major source of unnecessary complication is the conversion of variables during primitive recovery. origIGM uses the Noble2D primitive recovery routine, which it borrowed from the HARM code. Because of this, users will find three sets of variables in origIGM: its own, HARM’s, and those required by the Noble2D routine. The code converts variables from origIGM to HARM, and then from HARM to Noble2D, performing the opposite conversions after the conservative-to-primitive routine finishes. However, many of these conversions are effectively “do-nothing computations,” as origIGM and Noble2D share some variables that HARM does not. We have condensed this to a single variable conversion, removing several layers of interfaces in the process.
We have also cleaned up many auxiliary functions, such as the function that enforces the inequalities that $\tilde{\tau}$ and $\tilde{S}_{i}$ must satisfy in order to be physical. While not individually significant, these minor improvements simplify the logical flow of the code. As discussed in Appendix B, all of these changes are validated against the original code to ensure that their behavior is identical.
III.2 EoS Feature Extensions
We have also made several improvements and extensions to the latest version of IllinoisGRMHD in the Einstein Toolkit. The remaining subsections describe improvements to the evolution codes that implement GRHayL, and not the core library itself. This section details the extensions to IllinoisGRMHD’s EoS support. These include full support for hybrid polytropic EoSs, infrastructure to incorporate tabulated EoS and neutrino leakage into the official release version of origIGM, and the addition of the simple EoS.
origIGM originally only supported simple polytropic hybrid EoSs. While it had much of the code needed to extend to piecewise polytropic EoSs, elements of the code contained implicit assumptions regarding the EoS. Recent improvements have added tabulated EoS [Werneck_IGM] in a forked version of origIGM. With GRHayL, we fully incorporate the piecewise polytropic hybrid EoS feature into the release version of origIGM in the Einstein Toolkit, and we are working on tabulated EoS support. We also add the extra features necessary to use the simple EoS. While this is of less interest for physical simulations, the commonly used Balsara tests [Balsara_2001] assume this EoS, so the original origIGM code could not run those tests with the correct EoS for comparison with the exact solution.
III.3 More Boundary Condition Options
As a small addition, we provide a new boundary condition option. The hydrodynamic variables use copy boundary conditions by default, while outflow boundary conditions are applied to the velocities. However, the latter condition is incompatible with 1D or 2D tests, such as the Balsara tests. As such, we introduce the option to disable the outflow boundary condition to support 1D and 2D simulations.
III.4 Con2Prim Improvements
The most significant changes and improvements to pre-existing code are in the Con2Prim implementation. origIGM uses the Noble2D routine, with the Font1D routine as the only backup. While Font1D virtually guarantees a successful inversion, it does so by setting $P=P_{\rm cold}$, which introduces spurious cooling in the system. This means that any Con2Prim failures near the surface of the stars will cause discontinuities in the pressure due to the thermal component of the pressure suddenly disappearing wherever Font1D is applied, limiting the code’s accuracy.
To address these issues, IllinoisGRMHD supports up to four Con2Prim routines of the user’s choosing. This design allows users to employ any of GRHayL’s Con2Prim methods and control their ordering at runtime. For hybrid EoS with entropy advection evolution, a suggested combination is Noble2D, Palenzuela1D, Noble1D_entropy, and finally Palenzuela1D_entropy.
In IllinoisGRMHD, the following strategy is adopted by default in order to recover the primitive variables:
  1. 1.
    attempt each user-selected Con2Prim method once, in order, until one succeeds;
  2. 2.
    if all selected methods fail, attempt Font1D (for hybrid EoS only); and
  3. 3.
    if recovery still fails, perform an atmospheric reset.
This multi-stage approach is designed to improve the likelihood that Con2Prim returns valid primitive variables. It also avoids the use of Font1D and atmospheric reset for as long as possible so as to not discard any thermal contributions. Note that Font1D is currently not optional in IllinoisGRMHD, requiring the user to change the code in order to disable it.
III.5 Improved Stability and Accuracy
As we factored out algorithms from origIGM, we noticed a problematic pressure floor that is applied after Con2Prim routines, namely
$P\geq 0.9P_{\rm cold}\;.$
This floor was phenomenologically motivated, as choosing a value that is slightly less than $P_{\rm cold}$ allows for pressure overshoots at the neutron star surface, resulting in less diffusion and better preservation of central density. However, a more careful analysis reveals that flooring the pressure to a value less than $P_{\rm cold}$ leads to imaginary sound speeds. While such a choice of pressure floor may produce results that appear “better” from a dissipation perspective, it leads to clearly unphysical behavior and risks (very rare) failures due to division by zero. As such, we enforce the pressure floor $P\geq P_{\rm cold}$ instead.
Additionally, the computation of the source terms for the hydrodynamic variables has been improved. The original origIGM code first interpolates BSSN metric quantities to cell faces (which are needed for the flux calculations) using fourth-order interpolation. It then uses these face-centered values to compute the metric derivatives (used for the source terms) using a second-order finite difference. It is not immediately obvious what the order of accuracy of the result is, since the finite difference does depend on the larger stencil through the interpolated quantities. We modify GRHayLHD and IllinoisGRMHD to use fourth-order finite differences of cell-centered ADM quantities for the metric derivatives, which shows noticeable improvements in conserving the central star density. Section V.2 discusses the effect in more detail.
III.6 Memory Usage
While origIGM has proven to be a robust code, with other groups developing their own variants, the code does have some obvious places where it can be improved. One is the memory usage, which we have significantly reduced. We detail the changes below, but they ultimately result in a reduction in memory usage of over 40%. This percentage only refers to memory usage by the IllinoisGRMHD thorn, not the entire Einstein Toolkit, but it still represents a noticeable reduction of memory usage for simulations using IllinoisGRMHD.
The first major change is the removal of all BSSN grid variables from the origIGM thorn. The original code used both ADM and BSSN grid variables, often recomputing them in several places. We eliminated dependencies on BSSN grid variables, greatly simplifying the GRHayL code and removing non-essential grid variables from IllinoisGRMHD. We also improved the efficiency of the hydrodynamic RHS computations in GRHayL by generating optimized code via NRPy, which removed the need for additional grid variables to store $T^{\mu\nu}$.
Another change concerns how we compute reconstructed variables. Previously, all variables were reconstructed in separate loops, stored, and then used for the RHS calculations, even though most were never reused. In practice, only the reconstructed velocities are needed again, to compute doubly reconstructed velocities at cell edges for $A_{i}^{\mathrm{RHS}}$. Inside the IllinoisGRMHD thorn, we therefore combine reconstruction with the hydrodynamic-flux calculation and store only the reconstructed velocities.
All these changes result in a memory overhead reduction of over 40% from origIGM to IllinoisGRMHD. While entropy evolution and experimental tabulated EoS support do require a few additional grid variables, they still introduce far fewer variables than the number of variables that we have removed. The new GRHayLHD GRHD thorn has no magnetic variables, so it reduces the memory usage by $\sim$80% compared to origIGM. These changes represent a significant reduction of memory usage and access.
IV Code Infrastructures
IV.1 Einstein Toolkit
Before discussing our numerical results, we first describe in more detail the infrastructures in which we have implemented GRMHD codes using GRHayL. The Einstein Toolkit [ET_2023_11] is a collection of codes for astrophysics simulations. Most of these codes are built on the Cactus framework [Cactus], which provides a flexible infrastructure that allows many different code modules (called “thorns”) to be compiled together. The origIGM code is one such thorn. While Cactus provides many features, the actual numerical grids, parallelism, and related low-level details are implemented by driver thorns. The driver currently used for most production-level simulations is the Carpet driver, which provides AMR, implements MPI parallelism, and handles memory allocation during the simulation.
Similarly, CarpetX is a newly released driver that builds on the AMReX framework [amrex]. One major change with this driver is the automation of ghost-zone synchronization and boundary-condition application. This feature was added to Cactus in the PreSync update [Cupp_2018, Cupp_2019] but had not been implemented in many Carpet-based codes, including origIGM. CarpetX is designed with this new functionality in mind, and CarpetX-compatible codes are expected to provide the option for this more advanced inter-processor communication paradigm. Combined with the new looping and grid-function syntax, origIGM would need to be changed significantly to work with CarpetX. As such, an infrastructure-agnostic solution is highly desirable to reduce the work overhead of maintaining code for both drivers.
GRHayL provides this functionality within the Einstein Toolkit via the GRHayLib thorn, which directly compiles the GRHayL source code and sets up GRHayL parameters based on user inputs. This approach is more amenable to the Einstein Toolkit ecosystem than linking to an external library and improves ease of use within the Einstein Toolkit. Since GRHayLib only provides the core library functions, it is independent of the mesh driver.
We also introduce the GRHayLHD and IllinoisGRMHD thorns, which reimplement the features of the origIGM thorn using GRHayLib. GRHayLHD implements a pure GRHD code, and IllinoisGRMHD implements the GRMHD code. These thorns define grid variables, schedule functions with loops over the grid, and fill the data structures needed by the GRHayL functions. Everything else is external to these thorns. Thus, code duplication is minimized, and the pure hydrodynamic code benefits from the reduced computational cost and memory usage expected from removing the magnetic components.
Similarly, we provide a CarpetX-based thorn named GRHayLHDX, which again depends on GRHayLib. The IllinoisGRMHDX thorn is also in development, with expected inclusion in a future release of the Einstein Toolkit. With the inclusion of CarpetX in the Einstein Toolkit, providing reliable GRMHD codes for the new driver is critical for future science using the Einstein Toolkit, with AsterX [Kalinani:2024rbk] a noteworthy example. One important caveat is that CarpetX also supports GPUs. While work to support GPUs has begun, GRHayL currently only supports CPU code. As such, the CarpetX thorns run on the host at the moment. Once the library functions are available on the device, the thorns’ code can be easily switched to GPU-based computation.
IV.2 BlackHoles@Home
BlackHoles@Home is an infrastructure built on the code-generation package NRPy [Ruchlin:2018, Etienne_nrpytutorial]. NRPy allows for efficient generation of optimized C code from symbolic Python expressions. BlackHoles@Home leverages this Python package to produce minimal but efficient numerical relativity code using high-order finite-differencing and time-stepping algorithms with extrapolation and high-order radiation outer boundary conditions. The infrastructure takes advantage of NRPy’s C code generation to produce state-of-the-art code for binary black hole simulations. The core goal of the infrastructure is to reduce the memory usage and inherent inefficiencies of box-in-box AMR on Cartesian grids by using multi-coordinate, multi-patch grids.
Previously, this infrastructure only supported vacuum spacetime evolutions. We have developed two GRHayL-based GR(M)HD codes for BlackHoles@Home: the GRHD code GRoovy and its GRMHD extension NRPyCartMHD. GRoovy [Jacques:2024pxh] implements GRHayL within BlackHoles@Home, yielding a GRHD code for fully dynamical spacetimes that supports single-patch curvilinear coordinates via a reference-metric formulation [Montero:2012yr, Baumgarte:2012xy, Mewes:2020vic]. NRPyCartMHD is a specialization of GRoovy that extends it to GRMHD while restricting the coordinate system to Cartesian coordinates. We are currently developing a supercomputer-ready version of GRoovy that leverages Charm++ task-based parallelism through superB [Jadoo:2025lgm].
We note that the Einstein Toolkit thorns Baikal and BaikalVacuum for spacetime evolution are also NRPy-generated, and BlackHoles@Home uses the exact same core routines internally. This allows for more meaningful comparisons between the two infrastructures, since the spacetime evolution uses the same code.
V Code Tests
We approach testing and validation through two complementary avenues. First, the core GRHayL library has many function-level tests that are regularly performed. The details of our continuous integration testing methods are given in Appendix B. Second, we compare physics simulations produced in several different infrastructures with results from origIGM. Through this testing, we validate the behavior of the replacements for origIGM by showing that the differences between the GRHayL-based codes and origIGM are within expectations. Additionally, we demonstrate successful parallel development of GRMHD codes using GRHayL in BlackHoles@Home, Carpet, and CarpetX.
V.1 1D Shock Tests
Balsara et al. [Balsara_2001] provide a very challenging battery of tests using one-dimensional GRMHD flows characterized by sharp features—along with extreme Lorentz factors or magnetization—that push GRMHD codes to their limits. These tests also have analytic solutions, allowing for direct validation of a numerical code’s results. While origIGM is capable of passing these tests, one must remove its pressure floors and ceilings and adjust some EoS functions in order to do so, preventing them from being run without manual changes to the code. Due to our changes and improvements, IllinoisGRMHD can run these tests without modification by using the simple EoS option. Here we briefly describe the Balsara tests and present results comparing IllinoisGRMHD and GRHayLHD with analytical results and GRHydro, another trusted GRMHD code in the Einstein Toolkit.
Table 3: Basic setup for the left (right) side of the shock for the Balsara tests.
Test $t_{\mathrm{final}}$ $\Gamma$ $\rho$ $P$ $\tilde{u}^{x}$ $\tilde{u}^{y}$ $\tilde{u}^{z}$ $B^{x}$ $B^{y}$ $B^{z}$ $W$ $P/P_{\mathrm{mag}}$
[…]
2 $\frac{5}{3}$ 11 $30$ (11) 55 66 ($\frac{7}{10}$) 66 ($\frac{7}{10}$) $0.62$ ($0.077$)
3 $1000$ ($\frac{1}{10}$) $10$ 77 ($\frac{7}{10}$) 77 ($\frac{7}{10}$) $10.1$ ($0.002$)
[…]
5 $\frac{11}{20}$ $\frac{27}{25}$ (11) $\frac{19}{20}$ (11) $\frac{2}{5}$ ($\frac{9}{20}$) $\frac{3}{10}$ ($\frac{1}{5}$) $\frac{1}{5}$ 22 $\frac{3}{10}$ ($\frac{7}{10}$) $\frac{3}{10}$ ($\frac{1}{2}$) $1.19$ ($1.18$) $0.27$ ($0.34$)
The Balsara tests are all pure MHD tests in static Minkowski spacetime and have an initial shock placed at $x=0$. Note that the initial data for these tests prescribe the Valencia velocity $\tilde{u}^{i}$ instead of $v^{i}$, but in flat space $\tilde{u}^{i}=v^{i}$ (see Eq. 13). In Table 3 we display the initial parameters for the different Balsara tests, also listing the Lorentz factor and the ratio $P/P_{\rm mag}$.
As shown in the table, all of these tests have very small ratios of gas pressure to magnetic pressure. In addition, Balsara 3 has a pressure shock of four orders of magnitude without any change to the density, and the Balsara 4 test has an enormous Lorentz factor of over 22. Simply put, these tests all push an MHD code to the limits of the physics they are designed to model and serve as excellent first tests that an MHD code is working properly.
Figure 1: Comparison of GRHydro, IllinoisGRMHD, and GRHayLHDX with the Balsara 0 test.
Figure 2: Relative difference between IllinoisGRMHD and GRHayLHDX in the Balsara 0 test.
Figure 3: Comparison of GRHydro and IllinoisGRMHD against the exact solution for the Balsara tests. The left column shows the density at the final time of the test, and the right column shows the $y$ component of the magnetic field.
One drawback to the Balsara tests is that they all have non-zero magnetic fields, which means that the analytical results cannot be used for purely hydrodynamic codes. However, we can still use a modified version of the Balsara tests to compare different codes. Since the shocks are still significant even in just the hydrodynamic variables, this still serves as an excellent test of the purely hydrodynamic GRHayL codes. We therefore compare and validate the hydrodynamic codes using a “Balsara 0” test, obtained from Balsara 1 by setting the magnetic fields to zero.
In Fig. 1, we compare results from the Balsara 0 test with GRHydro, IllinoisGRMHD, and GRHayLHDX. All three codes show excellent agreement, validating the CarpetX-based code. Since IllinoisGRMHD and GRHayLHDX are both based on GRHayL, the only differences are that one explicitly evolves the magnetic variables and that they are implemented in different infrastructures. As such, this provides an excellent comparison between them.
In Fig. 2, we show the relative difference between the final results of the two GRHayL-based codes. The error is concentrated around the shocks or areas most affected by the shock propagation, as can be seen by examining the shock profile in Fig. 1 and comparing it with the errors in Fig. 2. The peak around zero represents lingering effects of the original shock, which started in the center.
As part of our validation, we also ran all five Balsara tests with IllinoisGRMHD and GRHydro, a well-known and tested code within the Einstein Toolkit. All the simulations have 1600 points in the shock direction and 8 points in the perpendicular directions. Except for the Balsara 3 test, the simulations all use PPM reconstruction and the HLL approximate Riemann solver for the hydrodynamic quantities. In the Balsara 3 test, GRHydro crashes due to Con2Prim failures when using PPM, so GRHydro instead uses minmod for the reconstruction. This failure mode is not unique to GRHydro, as Spritz also uses minmod for this test. We also see Con2Prim failures near the outer edges of the grid, but our backup routines prevent the run from ending prematurely. While GRHayL provides minmod, IllinoisGRMHD currently only supports PPM, so we cannot match GRHydro’s settings for this test.
The primary difference between the two codes is in the magnetic sector. GRHydro only supports evolving the magnetic field directly or evolving a cell-centered vector potential, while IllinoisGRMHD evolves a staggered (i.e. edge-centered) vector potential. We choose to compare against GRHydro with magnetic field evolution, which is the default and more heavily tested evolution scheme in the code. Further, it has been shown that evolution of the cell-centered vector potential leads to increased noise around shocks (see e.g. [Spritz]).
In addition to these numerical codes, we use the analytical solution produced by the code presented in [Giacomazzo_2006]. Fig. 3 shows the results for the Balsara tests with IllinoisGRMHD and GRHydro, overlaid with the exact solution. We see that the two codes agree very well not only with the exact solution, but also with each other. In the Balsara 3 test, IllinoisGRMHD captures the shock better due to using PPM instead of minmod.
V.2 TOV Neutron Star
To further validate GRHayL and its implementations within the Einstein Toolkit and the BlackHoles@Home infrastructures, we performed simulations with single neutron stars using Tolman–Oppenheimer–Volkoff (TOV) initial data. The TOV solution is a spherically symmetric solution to Einstein’s equations with a non-zero stress-energy tensor. In the following subsections we show numerical results from evolving TOV initial data in a fully dynamical, three-dimensional spacetime with no symmetry assumptions. All initial data used for these tests were generated using the TOV solver within NRPy.
V.2.1 Comparison across infrastructures
Our first suite of tests consists of evolving the same TOV initial data across different infrastructures, comparing the following codes: origIGM and IllinoisGRMHD within the Einstein Toolkit, and NRPyCartMHD. The initial central density of the star is $\rho_{c}=1.28\times 10^{-3}$ in code units, and we adopt an initially cold hybrid EoS with $K_{0}=100$ and $\Gamma_{0}=\Gamma_{\rm th}=2$. The TOV initial data do not include magnetic fields, making this a pure hydrodynamics test. The spacetime is evolved using Baikal [Ruchlin:2018, Etienne_nrpytutorial], available within both the Einstein Toolkit and BlackHoles@Home.
Baikal evolves the spacetime variables using the BSSN formulation [Brown_2009], and employs the $1+\log$ lapse [Campanelli_2006] and a second-order, non-covariant, advective shift evolution [Meter_2006] as gauge conditions, i.e.
$\partial_{t}\alpha$ $=\beta^{i}\partial_{i}\alpha-2\alpha K,$
$\partial_{t}\beta^{i}$ $=\beta^{j}\partial_{j}\beta^{i}+B^{i},$
$\partial_{t}B^{i}$ $=\beta^{j}\partial_{j}B^{i}+\frac{3}{4}\partial_{0}\bar{\Lambda}^{i}-\eta B^{i},$
where $K$ is the trace of the extrinsic curvature $K_{ij}$, $\bar{\Lambda}^{i}$ and $B^{i}$ are auxiliary variables, and $\eta$ is the shift damping parameter. We set $\eta=1/3$, use Kreiss–Oliger dissipation at a strength of $0.2$ [Kreiss_1973], and use fourth-order finite differencing for the spacetime fields. For the hydrodynamic fields we use PPM reconstruction, the Noble2D conservative-to-primitive solver, and set the atmospheric density to $\rho_{\mathrm{atm}}=\rho_{c}\times 10^{-9}$. In all tests we ensure that the diameter of the star is sampled with at least ${\approx}30$ points along a coordinate axis. These codes evolve the initial data forward in time using the method of lines and an RK4 scheme and, apart from NRPyCartMHD, use five levels of fixed mesh refinement (FMR) in Cartesian coordinates with the outer boundary placed at ${\pm}240$ in all three directions. We evolve the initial data at two resolutions, MR and HR, corresponding to resolutions on the coarsest level of $dx=8.0$ and $dx=5.0$, respectively. Since NRPyCartMHD supports only a single Cartesian grid, we place its outer boundary at ${\pm}40$ in all three directions and choose the grid spacing to match that of the finest refinement level of the FMR hierarchy.
In the top panel of Fig. 4 we show results of the time evolution of the normalized change in central density for all codes, truncating the results from NRPyCartMHD when outer boundary effects become apparent, with all the GRHayL-based codes using the full suite of changes from Sec. III. These results demonstrate that the frequencies of the fundamental modes of oscillation for the star are recovered [Font_2000_nonlinear_tests, Font_2002]. Further, the amplitude of these oscillations, which are driven by discretization errors in the nominally static initial data, is expected to converge to zero with increasing resolution. In the bottom panel of Fig. 4 we plot the approximate convergence order of the error in central density for origIGM and IllinoisGRMHD. We observe that they have a similar convergence order at around $n=2.5$, consistent with what was found in previous studies (e.g., [Werneck_IGM]).
Figure 4: Top: Comparison of non-magnetized TOV initial data evolution between origIGM, IllinoisGRMHD, GRHayLHDX, and NRPyCartMHD, each at two resolutions. IllinoisGRMHD and NRPyCartMHD are equivalent until outer boundary effects are apparent, therefore we truncate the results of NRPyCartMHD before they manifest. Bottom: Comparison of convergence order between origIGM and IllinoisGRMHD. IllinoisGRMHD consistently outperforms origIGM.
Figure 5: Comparison of non-magnetized TOV initial data evolution between origIGM and various changes made to IllinoisGRMHD during development.
In Fig. 5 we plot the normalized change in central density over time, comparing origIGM and different versions of IllinoisGRMHD. These versions represent the changes described in Sec. III. The second-order data best match the origIGM code, and the central density has similar behavior and loss due to diffusion. As discussed in Sec. III.2, the local grid data needed to perform fourth-order finite differencing are already available, so we can immediately change the derivative calculation without requiring any extra data. This change better preserves the central density.
On the other hand, allowing the pressure to fall below the cold pressure leads to cooling and limits diffusion. Therefore, the change to the pressure floor results in more diffusion. As discussed in Sec. III.5, increased diffusion is a necessary sacrifice to avoid unphysical behavior and possible divisions by zero.
Figure 6: Relative difference between IllinoisGRMHD and NRPyCartMHD, comparing two outer boundary locations while keeping the resolution fixed and using a single refinement level.
Finally, to confirm the effective integration of GRHayL within both the Einstein Toolkit and BlackHoles@Home infrastructures, we plot the relative difference between the two in Fig. 6, using a single Cartesian mesh for both. We observe that as we vary the location of the outer boundary, while keeping the resolution fixed, the duration over which we maintain round-off-level agreement is extended. This demonstrates that the observed discrepancies arise solely from the slightly different implementations of Sommerfeld radiation boundary conditions in Carpet and BlackHoles@Home.
V.3 Piecewise polytrope EoS study
[…]
Figure 7: Same as Fig. 4 but for piecewise polytropic SLy EoS.
Figure 8: Fourier transform of time series of the central density evolution for a TOV in dynamical spacetime with the piecewise polytropic representation of the SLy EoS. Vertical dashed lines indicate the frequencies of the lowest four radial oscillation modes for the SLy EoS [Sen:2022kva].
V.4 BNS Inspiral, Merger, and Collapse of HMNS
As a final test, we consider BNS mergers, which pose a particularly challenging problem for any GRMHD code. At present, however, GRoovy is implemented only in the single-curvilinear-patch version of BlackHoles@Home. Although multipatch support is under development, it is not practical to perform BNS simulations within a single grid patch. The current development status of CarpetX similarly restricts the simulations we can perform: several key capabilities, including BNS initial data and various GRHD diagnostics, are not yet available in the new driver. Given these limitations and the high computational cost of BNS simulations, in this work we restrict our attention to comparisons between origIGM and IllinoisGRMHD using Carpet.
V.4.1 Hybrid EoS, Simple Polytrope
Figure 9: Comparison of the maximum density for origIGM and IllinoisGRMHD during inspiral with both perturbed and unperturbed simulations. The $y$-axis is rescaled by the maximum density at $t=0$.
Figure 10: Comparison of the maximum density for origIGM with randomly perturbed initial data from hypermassive neutron star formation to collapse.
For the hybrid EoS comparison, we use $\Gamma=\Gamma_{\mathrm{th}}=2$ and $K\approx 123.64$. The limits on the density are set to $\rho_{\mathrm{min}}=\rho_{\mathrm{atm}}=10^{-13}$ and $\rho_{\mathrm{max}}=0.003=3.1\rho_{c,0}$, where $\rho_{c,0}$ is the initial central density of the neutron stars. The primary Con2Prim method is set to Noble2D, with backups of Palenzuela1D, Noble1D_entropy, and finally Font1D. This does differ from origIGM, which uses the only available backup of Font1D.
We use LORENE initial data for an equal-mass BNS system with an initial separation of $45$ km. Each star has mass $M\approx 1.625\,M_{\odot}$ and radius $R\approx 13.67\,\mathrm{km}$. We use the Seed_Magnetic_Fields thorn to set up the initial magnetic fields and the Baikal thorn to evolve the spacetime. Further details regarding the AMR grids and thorn setup are provided in the hybrid BNS parameter file packaged with the IllinoisGRMHD thorn.
Due to the various improvements (detailed in Sec. III), the codes no longer agree at the round-off-level. To assess their level of agreement, we perform three BNS simulations with both evolution codes to analyze their behavior. One is the control with no perturbations, and the other two perturb the initial $\bm{P}$ data at the 9th and 14th significant digits, respectively.
Fig. 9 shows the quantity $\rho_{\mathrm{max}}/\rho_{0}$ during the inspiral phase of the simulation for all six simulations. The two codes show the same behavior, and the differences follow the same trends as the TOV simulations. At this stage of the simulation, the perturbations show no significant effect on the central density of the stars. The central densities dissipate away faster in IllinoisGRMHD due to the change to the $P_{\mathrm{min}}$ floor, which has a slightly stronger effect in these simulations than in the TOV simulations.
The perturbative effects start significantly affecting the evolution during the hypermassive neutron star phase, where even small changes can affect the evolution of the system. Fig. 10 compares $\rho_{\mathrm{max}}$ from the origIGM simulations (unperturbed and perturbed at the 14th digit) with that from the unperturbed IllinoisGRMHD simulation. We cut off the data once the simulation reaches the maximum allowed density for the simulation. As shown, perturbations to initial data cause the number of oscillations to change in the original origIGM code. The new IllinoisGRMHD has round-off-level differences at every Con2Prim evaluation relative to the original code due to code improvements and changes, which would similarly cause differences between the old and new results. Despite these cumulative perturbations, IllinoisGRMHD’s results are still on the order of a round-off-level perturbation relative to the original code. Given the unstable nature of the hypermassive neutron star and the repeated perturbations caused by various code improvements and simplifications, IllinoisGRMHD displays remarkable agreement with origIGM.
origIGM IllinoisGRMHD
(a) $t=0$.
(b) $t=0$.
(c) $t=944$ (two full orbits).
(d) $t=944$ (two full orbits).
(e) $t=1632$ (first touch).
(f) $t=1632$ (first touch).
Figure 11: Log plots of data from a magnetized, equal-mass BNS during inspiral with a hybrid EoS. The left side shows the xyxy-plane for origIGM, and the right side shows the same for IllinoisGRMHD. For all panels, the right half shows the density, and the left half shows the magnetic field strength $\sqrt{b^{\mu}b_{\mu}}$.
origIGM IllinoisGRMHD
(a) $t=1920$ (hypermassive neutron star).
(b) $t=1920$ (hypermassive neutron star).
(c) $t=2768$ (half-way to collapse).
(d) $t=2768$.
(e) $t=3600$ (post-collapse).
(f) $t=4640$ (post-collapse).
Figure 12: Same as Fig. 11 for the post-merger phase.
Figure 13: Log plots of density (left) and magnetic field strength (right) at $t=1216$. Both plots are split with data from IllinoisGRMHD (left) and origIGM (right).
Figure 14: Log plots of density (left) and magnetic field strength (right) at $t=1344$. Both plots are split with data from IllinoisGRMHD (bottom) and origIGM (top).
Of course, simply tracking the central density does not fully assess the physics of the simulation. To further examine the results, we show spatial slices of density and magnetic field strength $\sqrt{b^{\mu}b_{\mu}}$ for both codes. In Fig. 11, we show the inspiral at three different times. As expected, the initial data in Fig. 11a and Fig. 11b match exactly. While it is not as clear with still images, the distortions that appear in the magnetic field strength during the inspiral are noticeably worse with origIGM. The stellar surface in Fig. 11c and Fig. 11e shows far worse behavior than in Fig. 11d and Fig. 11f. The chipped appearance of the isocontours is completely absent in the new code. While it develops some asymmetric aspects during the inspiral, the magnetic fields are advected better than in the original code.
In Fig. 12, we show three spatial slices for the post-merger. Note that the collapse time is different between the codes. The first two sets of plots are for the same time and display very similar behavior. The magnetic field strength again follows the matter flows better in IllinoisGRMHD, with the matter tails coming off the merging binary having accompanying magnetic components that are missing in origIGM. Fig. 12e and Fig. 12f show the post-collapse system after the central mass has settled down into a black hole with inspiraling matter. Both codes show very similar dynamics and remain stable out to $t=6000$.
To illustrate the differences more clearly, Fig. 13 and Fig. 14 show the density and magnetic field strength for both codes. origIGM has very noticeable distortions in magnetic field topology near the star surface, while IllinoisGRMHD has a much smoother profile. Further, the magnetic field in IllinoisGRMHD follows the distortion of the stars, with the deformation leading to merger causing a deformation of the magnetic field strength, as seen in Fig. 13. In contrast, origIGM remains nearly spherical.
As the stars enter the merger, the inspiral rate outpaces the magnetic field’s ability to keep up, leading to the bulge in magnetic field strength visible in Fig. 14. The deformation towards positive $x$ for the lower star is the same deformation present in Fig. 13, as this feature fails to follow the star’s rotation through the last quarter-rotation into the merger. This is not apparent in origIGM since it remains nearly spherical all the way to merger.
V.4.2 Hybrid EoS, Piecewise Polytrope
[…]
We set up a grid with eight refinement levels by factors of two, with the resolution at the finest level $\Delta x_{8}\approx 185\penalty 10000\ \mathrm{m}$. After BH formation, two additional refinement levels are added to resolve the puncture, making the highest grid resolution $\Delta x_{10}\approx 46\penalty 10000\ \mathrm{m}$.
Figure 15: Relative error in maximum density over time between IllinoisGRMHD and origIGM for BNS simulations with (blue, solid) and without (orange, dashed) magnetic fields.
[…]
origIGM IllinoisGRMHD
(a) $t=0$.
(b) $t=0$.
[…]
VI Concluding remarks
GRHayL provides an effective method for quickly developing GRMHD evolution code in new infrastructures. Any changes to the library are validated by continuous integration testing (see Appendix B). Few GRMHD codes perform regular testing at the level of individual functions (and demand round-off-level agreement), and these tests represent professional industry standards that are rarely implemented in open-source scientific codes.
The current library contains all the necessary pieces to quickly implement an IllinoisGRMHD-like code in new infrastructures. This includes multiple EoS types, Con2Prim and reconstruction methods, hydrodynamic and induction equation RHSs, and neutrino leakage. In addition to the current features, plans include adding more advanced radiation transport schemes, more atmosphere prescriptions and more Con2Prim options in the near future. Work is already underway to extend the Noble implementations to support tabulated EoSs, as well as adding the 2D and 3D Cerdá-Durán et al. routines [Cerda_2008]. Support for M1 closure neutrino transport based on [Radice:2021jtw] is also planned.
Implementation of BlackHoles@Home’s multi-patch, multi-coordinate infrastructure within GRoovy will enable highly efficient BNS simulations. Work also continues towards adding magnetic evolution to the curvilinear code, which must be handled with particular care due to the combination of staggered grids and coordinate singularities.
Several more GRMHD implementations using GRHayL are already in the planning or development phase for other infrastructures, some of which currently lack hydrodynamics codes. While the community benefits from multiple codes for validation and scientific rigor, the advantages of having a code that can easily be integrated into any infrastructure are significant. These advantages include easier vetting of new infrastructures, faster implementation of GRMHD code in future HPC environments, and establishing a cross-infrastructure “standard” that can more easily determine whether differences between codes arise from the physics code or the underlying infrastructure.
Acknowledgements
We thank Erik Schnetter and Roland Haas for their help with using the new CarpetX driver. We also thank Jay Kalinani for discussions about the AsterX GRMHD code which also uses CarpetX, as well as discussions regarding implementing the simple EoS.
T. Pierre Jacques gratefully acknowledges funding support from NASA FINESST-80NSSC23K1437, West Virginia University’s Chancellor’s Scholarship program, and the Southern Regional Education Board’s Dissertation Award program. Z. B. Etienne gratefully acknowledges funding support from NSF awards AST-2227080, PHY-2110352, OAC-2227105, and NASA awards ISFM-80NSSC21K1179 and TCAN-80NSSC18K1488.
This research made use of the resources of the High Performance Computing Center at Idaho National Laboratory, which is supported by the Office of Nuclear Energy of the U.S. Department of Energy and the Nuclear Science User Facilities under Contract No. DE-AC07-05ID14517.
Appendix A Basic Equations
For completeness, we provide a brief overview of the GRMHD equations and choice of variables used by GRHayL and the core equations that it is designed to solve. As is standard in a numerical relativity framework, we use the 3+1 formalism, in which the spacetime metric is written as
$ds^{2}=\left(-\alpha^{2}+\beta^{i}\beta_{i}\right)dt^{2}+2\beta_{i}\,dt\,dx^{i}+\gamma_{ij}\,dx^{i}\,dx^{j}\;,$ (6)
where $\alpha$ is the lapse function, $\beta^{i}$ is the shift vector, and $\gamma_{ij}$ is the ADM physical spatial metric.
The GRMHD equations—conservation of baryon number, conservation of lepton number, conservation of energy-momentum, and homogeneous Maxwell’s equations—are given by
$\nabla_{\mu}\left(n_{\mathrm{b}}u^{\mu}\right)$ $=0\;,$ (7)
$\nabla_{\mu}\left(n_{\mathrm{e}}u^{\mu}\right)$ $=0\;,$ (8)
$\nabla_{\mu}T^{\mu\nu}$ $=0\;,$ (9)
$\nabla_{\mu}{}^{*}\!F^{\mu\nu}$ $=0\;,$ (10)
respectively. Here, $n_{\mathrm{b}}$ and $n_{\mathrm{e}}$ are the number densities of baryons and leptons, $\rho=n_{\mathrm{b}}m_{\mathrm{b}}$ the baryon density, $m_{\mathrm{b}}$ the baryon mass, and $Y_{\mathrm{e}}=n_{\mathrm{e}}/n_{\mathrm{b}}$ the electron fraction. Further, $u^{\mu}$ is the fluid four-velocity, $^{*}\!F^{\mu\nu}=(1/2)\tilde{\epsilon}^{\mu\nu\rho\sigma}F_{\rho\sigma}$ is the dual of the Faraday tensor $F^{\mu\nu}$, and $\tilde{\epsilon}^{\mu\nu\rho\sigma}$ is the Levi-Civita tensor. Finally, we assume ideal MHD ($u_{\mu}F^{\mu\nu}=0$) throughout.
The evolution equations are written in flux-conservative form via
$\partial_{t}\bm{C}+\nabla_{j}\bm{F}^{j}=\bm{S}\;,$ (11)
where $\bm{C}$ is the vector of conservative variables, $\bm{F}^{j}$ is the flux vector along direction $j$, and $\bm{S}$ is the vector of source terms. The vectors $\bm{C}$, $\bm{F}^{j}$, and $\bm{S}$ depend directly on the so-called primitive variables $\bm{P}$, defined as
$\bm{P}=\begin{bmatrix}\rho\\ Y_{\mathrm{e}}\\ S\\ P\\ v^{i}\\ B^{i}\end{bmatrix}\;$. (12)
where $P$ is the pressure, $v^{i}\equiv u^{i}/u^{0}$ is the fluid three-velocity, and $S$ is the entropy. Note that this velocity is different from the velocity used in the Valencia formulation [Banyuls_1997, Marti_1991]
$\tilde{u}^{i}=\alpha^{-1}\bigl(v^{i}+\beta^{i}\bigr)\;,$ (13)
which is measured by Eulerian observers with unit four-velocity $n^{\mu}=\left(\alpha^{-1},\alpha^{-1}\beta^{i}\right)$ normal to the spatial hypersurfaces.
Similarly, there are two common definitions of $B^{\mu}$, the magnetic field measured by Eulerian observers. The first, adopted by e.g., [Duez_IGM], is
$B^{\mu}=n_{\nu}{}^{*}\!F^{\mu\nu}\;,$ (14)
while the second, adopted by e.g., [GRHydro], is
$B^{\mu}=\frac{1}{\sqrt{4\pi}}n_{\nu}{}^{*}\!F^{\mu\nu}\;.$ (15)
Unlike the original version of IllinoisGRMHD, which adopts Eq. (14), GRHayL adopts Eq. (15). This choice does not affect the evolution scheme in any way, but it is the more natural choice from a numerical perspective, avoiding unnecessary multiplications and divisions by factors of $\sqrt{4\pi}$.
The conservative variables $\bm{C}$ (minus the magnetic field) in Eq. (11) are given by
$\bm{C}=\begin{bmatrix}\rho_{*}\\ \tilde{Y}_{\mathrm{e}}\\ \tilde{S}\\ \tilde{\tau}\\ \tilde{S}_{i}\end{bmatrix}\equiv\sqrt{\gamma}\begin{bmatrix}W\rho\\ W\rho Y_{\mathrm{e}}\\ WS\\ \tau\\ S_{i}\end{bmatrix}=\sqrt{\gamma}\begin{bmatrix}W\rho\\ W\rho Y_{\mathrm{e}}\\ WS\\ \alpha^{2}T^{00}-W\rho\\ \alpha T^{0}_{\ i}\end{bmatrix}\;$, (16)
where $\gamma=\det\left(\gamma_{ij}\right)$ and $W=\alpha u^{0}$ is the Lorentz factor. Note that $\rho_{*}$ is sometimes referred to in the literature as $\tilde{D}$. The adopted stress-energy tensor is that of a perfect fluid
$T^{\mu\nu}=\left(\rho h+b^{2}\right)u^{\mu}u^{\nu}+\left(P+P_{\mathrm{mag}}\right)g^{\mu\nu}-b^{\mu}b^{\nu}\;,$ (17)
where $h=1+\epsilon+P/\rho$ is the specific enthalpy, $\epsilon$ is the specific internal energy, $g_{\mu\nu}$ is the four-metric, $P_{\mathrm{mag}}=b^{2}/2$ is the magnetic pressure, $b^{2}\equiv g_{\mu\nu}b^{\mu}b^{\nu}$ is the magnetic energy density, and $b^{\mu}$ is the magnetic field measured by an observer comoving with the fluid (i.e., with four-velocity $u^{\mu}$),
$b^{0}$ $=u_{i}B^{i}/\alpha\;,$ (18)
$b^{i}$ $=\left(B^{i}/\alpha+b^{0}u^{i}\right)/u^{0}=\left(B^{i}+B^{j}u_{j}u^{i}\right)/W\;.$ (19)
To complete our prescription of Eq. (11), the flux terms may be written as
$\bm{F}^{j}=\begin{bmatrix}\rho_{*}v^{j}\\ \tilde{Y}_{\mathrm{e}}v^{j}\\ \tilde{S}v^{j}\\ \alpha^{2}\sqrt{\gamma}T^{0j}-\rho_{*}v^{j}\\ \alpha\sqrt{\gamma}T^{j}_{\ i}\end{bmatrix}\;$, (20)
and the source terms as
$\bm{S}=\begin{bmatrix}0\\ 0\\ 0\\ \alpha\sqrt{\gamma}\left[\Theta^{kl}K_{kl}-\left(T^{00}\beta^{k}+T^{0k}\right)\partial_{k}\alpha\right]\\ \frac{1}{2}\alpha\sqrt{\gamma}T^{\mu\nu}g_{\mu\nu,i}\end{bmatrix}\;$, (21)
where $\Theta^{kl}\equiv T^{00}\beta^{k}\beta^{l}+2T^{0k}\beta^{l}+T^{kl}$.
Appendix B Continuous Integration Testing
The industry standard for code testing is to have regularly scheduled tests which validate individual components, or units. These unit tests validate each individual component so that if a change breaks the code, it is clear which piece has diverged from the expected behavior. Since many codebases are worked on by large groups of people who may be implementing many different features simultaneously, it is standard to use a technique known as continuous integration (CI) testing. CI runs unit tests every time the triggers are met, such as whenever someone pushes to the repository. The library is also indirectly tested via the Einstein Toolkit’s CI, as several tests now include GRHayL-based thorns. However, we only discuss testing performed by the library and not third parties.
We perform CI testing via GitHub Actions. We trigger these tests every time a push is made to the main branch or when a pull request is opened or updated. A separate repository (GRHayL_TestData) contains all the data needed to run the tests. The CI tests pull the individual data files they need using curl, run the GRHayL functions using the initial data, and then validate the output data against a trusted version from origIGM whenever possible. These tests are run on a variety of operating systems, versions, and compilers.
This method of automated unit testing is a required feature of most professional codes, and adopting these standards ensures that our code library continues to meet our expectations of accuracy and compatibility across systems, compilers, and infrastructures. As new code is introduced to GRHayL, the list of unit tests is expanded to validate the additions, ensuring that all the components of the library can be trusted to have identical behavior in the future (barring purposeful improvements or bug fixes which cause changes to behavior).
B.1 Code Test Methodology
GRHayL has been extensively tested, and automated tests immediately validate any changes to the code. The tests read input data $u_{\mathrm{in}}$ and compute an output $v_{\mathrm{new}}$. We compare this against a trusted reference output $v_{\mathrm{old}}$. To estimate round-off-level sensitivity, we additionally compute $v_{\mathrm{pert}}$ by perturbing $u_{\mathrm{in}}$ and re-running the trusted code.
We define the error bars as
$\left|\frac{v_{\mathrm{new}}-v_{\mathrm{old}}}{v_{\mathrm{old}}}\right|<4\left|\frac{v_{\mathrm{pert}}-v_{\mathrm{old}}}{v_{\mathrm{old}}}\right|\;.$ (22)
The multiplication by 4 approximates the fact that the perturbed errors are unlikely to align to the maximum error. This error checking is very strict, far more than most codes demand for changes. Many codes set somewhat arbitrary error bars by using explicit relative and absolute tolerances that they hope correspond to the expected (worst-case) numerical error of the test. By estimating the numerical round-off error, we can more confidently state that the functions in GRHayL produce identical results to those produced by previous versions of the code.
As previously mentioned, the data used for these comparisons is kept in the GRHayL_TestData repository. This serves to keep GRHayL as small as possible and also permits more and larger tests since we do not have to be as worried about the file sizes. These tests are optionally compiled along with the library, and the source code is in the Unit_Tests directory. Any test that requires data from the GRHayL_TestData repo also has a data generation file in the data_gen subdirectory.
The tests can be separated into two categories. The tests prefaced with ET_Legacy have matching code in origIGM. As such, the only data generated by GRHayL is the input data. A separate branch of origIGM (GRHayLTestPatch) contains the code changes and a parfile to read in this data. The branched origIGM code produces the trusted outputs $v_{\mathrm{old}}$ and $v_{\mathrm{pert}}$ as binary files for all the ET_Legacy tests (which are also stored in the GRHayL_TestData repository). The GRHayL library computes $v_{\mathrm{new}}$ from the same input data, and these values are compared to $v_{\mathrm{old}}$ and $v_{\mathrm{pert}}$, ensuring that the GRHayL code matches the progenitor code in origIGM. For new code, the data and data-generating code are grouped into directories by gem. The data-generating code for functions new to GRHayL produces input, output, and perturbed output binary files.
All tests (with one exception) perturb the input data with
$u_{\mathrm{pert}}=u_{\mathrm{in}}\Bigl[1+\mathrm{rand}\bigl(-10^{-14},10^{-14}\bigr)\Bigr]\;.$ (23)
By choosing such a small perturbation we demand a very high level of agreement across code versions, compilers, and operating systems. The benefit of using perturbed values as error bars is that we need not set a single tolerance cutoff for the test. Each data point in the test effectively has its own customized cutoff value. The common implementation of tolerance cutoffs requires a global tolerance value. This can, at best, match the error of the worst part of the variable space. Areas where the code should have greater accuracy can therefore be assessed, rather than sliding under the limit set by less accurate parts of the variable domain.
As mentioned, one test uses different perturbed input values. The comparison between the origIGM and GRHayL Con2Prim routines requires some relaxation of the stringent requirements on GRHayL tests. This is because the differences between the two are more significant than those found in other pieces of origIGM. Although origIGM is a long-standing and well-tested code, many aspects of the Con2Prim routine allow for extensive simplifications and reductions in computation. This results in many (accumulating) round-off-level differences before even reaching the Con2Prim routine itself. These effective perturbations are larger than the “perturbed” input data that we are using to approximate numerical error. These perturbations are further magnified by going through the Newton-Raphson method of Noble2D and several more variable conversions before returning to the top-level code. As such, we relax the error bars by changing the perturbation factor from $10^{-14}$ to $10^{-12}$ in Eq. (23). Another problem with this comparison is that the Noble2D Con2Prim routine used by both can have catastrophic cancellation under the right circumstances. This affects the computed pressure and can lead to apparent code failures. This only occurs for very small pressures, so this one test has, in addition to larger perturbed inputs (effectively a larger relative tolerance value), a larger absolute tolerance cutoff for the pressure.
B.2 Code Coverage Tracking
Another concern for high-quality code testing is ensuring that the code is truly being exercised. If a test is set up incorrectly but passes, it may not be obvious that the code is not, in fact, being tested. To remedy this, we use gcovr (the recursive version of gcov) to track which lines of code are covered by our tests. Since we have many tests covering the library, compiling a single report from all the individual gcovr reports would be quite tedious. We instead link our CI tests to Codecov, which collates all the data and provides a single report on code coverage, including the percentage of lines covered. This also includes coverage percentage information for subdirectories. At the level of individual files, gcovr even reports line-by-line code coverage information. This information is very helpful for designing future tests and making sure edge cases are being properly tested.